# VisioPhysioENet: Multimodal Engagement Detection using Visual and Physiological Signals

### v1

@@ III-C Multimodal Fusion Phase
In our proposed architecture, we implement a Late Fusion strategy to effectively integrate Visual and Physiological features for enhanced classification performance. The visual features extracted in section III-A £$\pounds_{2}^{V}$ are processed using AdaBoost Classifier, to compute the probability distribution, $P_{AB}(V)$.
@@
On the other hand, the Physiological features from £$\pounds_{2}^{P}$ computed in section III-B are used for assessing the subject’s physiological responses using a Random Forest Classifier.
@@
In the final stage, a Late Fusion technique using stacked generalization combines the probabilities $P_{AB}$ and $P_{RF}$ using a meta-classifier (logistic regression), producing the final classification.
@@ IV-A1 Dataset and Implementation
The proposed system was evaluated using the train and test subsets from the DaiSEE dataset [27], which includes 9,068 videos from 112 adults, annotated by human experts. The model training used an 80-20 split and 10-fold cross-validation, conducted on an Intel(R) Core(TM) i7-7700, 3.70 GHz and Nvidia RTX 2070 GPU.
@@ IV-A2 Models
We developed three baseline models to determine the proposed system’s architecture, which is also guided by the ablation studies in Section IV-D. Baseline 1 utilized visual features such as eye category, head position, and eye gaze direction; Baseline 2 employed rPPG physiological signals; and Baseline 3 integrated both modalities via early fusion whereas the proposed system integrates them using late fusion.
@@ IV-B Quantitative Results
The performance of our proposed method is detailed in Table I, where it outperforms state-of-the-art approaches, including the approach by Vedernikov et al. [28] that utilizes both visual and physiological modalities.
TABLE I: Results’ comparison with state-of-the-art. Here, ‘V,’ ‘P’ and ‘Acc’ denote visual & physiological modalities and accuracy; best & second best results are bold and underlined.
@@
[…]
@@ IV-C Qualitative Results
Fig. 2 shows sample results. Visual metric head position is influenced by pitch, yaw, and roll; eye gaze direction from landmarks and eye openness category from EAR. Physiological metrics, derived from the BVP signal, include HR and P2P interval, where crests and troughs indicate systolic and diastolic peaks.
@@ IV-D1 Choice of Appropriate Fusion Approach
Table II showcases our fusion approach’s impact. Stacked generalization incorporates error correction from both AdaBoost and Random Forest while enhancing robustness through logistic regression as the meta-classifier. The late fusion improves engagement prediction accuracy to 63.09%, outperforming all baselines.
@@
[…]
IV-D2 Computational requirements analysis
Sequential task execution leads to idle processor time. To improve efficiency, we used ThreadPoolExecutor, allowing tasks to run in parallel across multiple threads while optimizing CPU usage. Table III highlights the performance gains from parallel processing compared to sequential methods.
TABLE III: Comparison of Sequential (ST) and ThreadPoolExecutor (TPE) times in seconds.
@@ IV-D2 Computational requirements analysis
V Conclusion
In this work, VisioPhysioENet, a multimodal system combining visual cues and physiological signals to detect learner engagement, has been introduced. With its lightweight design and real-time processing, VisioPhysioENet achieved 63.09% accuracy on the DAiSEE dataset, surpassing state-of-the-art methods. It balances accuracy with efficiency, using feature estimation methods provided by Dlib and OpenCV for visual features, and the POS method for physiological signals, making it ideal for real-time use. However, the system is sensitive to environmental changes and needs more testing in natural settings. Future work will enhance scalability, add modalities, and refine it for varied educational contexts, focusing on real-world adaptability.

### v2

@@ III-C Multimodal Fusion Phase
In our proposed architecture, we implement a Late Fusion strategy to effectively integrate Visual and Physiological features for enhanced classification performance. The visual features extracted in section III-A £$\pounds_{2}^{V}$ are processed using AdaBoost Classifier [38], to compute the probability distribution, $P_{AB}(V)$.
@@
where $P_{\text{AB}}(V)$ is the probability distribution output from the AdaBoost classifier based on the visual features.
On the other hand, the Physiological features from £$\pounds_{2}^{P}$ computed in section III-B are used for assessing the subject’s physiological responses using a Random Forest Classifier [39].
@@
where $P_{\text{RF}}(P)$ is the probability distribution output from the Random Forest classifier based on the physiological features.
In the final stage, a Late Fusion technique using stacked generalization [40] combines the probabilities $P_{AB}$ and $P_{RF}$ using a meta-classifier (logistic regression [41]), producing the final classification.
@@
This Late Fusion approach allows our model to capitalize on the complementary strengths of AdaBoost and Random Forest, ensuring that both visual and physiological information is utilized to its full potential. The result is a more accurate and robust classification model that effectively synthesizes multi-modal data to make well-informed predictions.
@@ IV-A1 Dataset and Implementation
The proposed system was evaluated using the train and test subsets from the DAiSEE dataset [15], which includes 9,068 videos from 112 adults, annotated by human experts. The model training used an 80-20 train-test split and 10-fold cross-validation, conducted on an Intel(R) Core(TM) i7-7700, 3.70 GHz and Nvidia RTX 2070 GPU.
@@ IV-A2 Models
To establish a comprehensive understanding of the proposed system’s efficiency, three distinct baseline models were developed based on the insights from ablation studies discussed in Section IV-D.
  • Baseline 1: Utilizes visual features, including eye category, head positioning, and gaze direction, to analyze visual cues from the data.
  • Baseline 2: Processes physiological signals such as heart rate and the cumulative sum of systolic and diastolic peaks, focusing on the physical aspects of the dataset.
  • Baseline 3: Combines both visual and physiological modalities through an early fusion technique, serving as a foundational comparison for the proposed system which applies a more sophisticated late fusion approach for modality integration.
@@ IV-B Quantitative Results
Our system’s performance, as shown in Table I, underscores its superiority over existing approaches by incorporating a novel integration of visual and physiological data. Notably, it outperforms the model by Vedernikov et al. [42], which also combines these modalities but with less efficacy. The proposed system achieves an accuracy improvement of 8.6% over this model, demonstrating the effectiveness of the late fusion approach in handling multimodal data. This significant improvement in performance highlights the potential of our methodology to advance research in the field of multimodal systems, setting a new benchmark for future studies.
TABLE I: Results’ comparison with state-of-the-art. Here, ‘V,’ ‘P’ and ‘Acc’ denote visual & physiological modalities and accuracy
@@
VisioPhysioENet V+P 63.09%
ResNet + TCN[6] V 63.90%
Ordinal TCN[44] V 67.40%
Despite not achieving the highest accuracy, VisioPhysioENet stands out by leveraging both physiological and visual data, providing a richer, more nuanced understanding of engagement. It outperforms the only other multimodal model by 8.6%, demonstrating its superior ability to integrate these data types. The unique features of VisioPhysioENet, such as advanced feature extraction and fusion techniques, enhance its utility in real-world educational settings. Additionally, the robustness and generalizability of the model suggest it can effectively handle variations and imbalances across different datasets, making it a valuable tool for practical applications.
@@ IV-C Qualitative Results
Fig. 2 shows sample results. Visual metrics like head position is influenced by pitch, yaw, and roll; eye gaze direction by landmarks while eye openness category from Eye-Aspect Ratio (EAR). Physiological metrics, derived from the BVP signal, include heart-rate (HR) and peak-to-peak (P2P) interval, where crests and troughs indicate systolic and diastolic peaks, respectively.
@@ IV-D1 Choice of Appropriate Fusion Approach
The following studies have been carried out to select the appropriate architecture for the proposed model.
Table II presents our fusion approaches, highlighting how stacked generalization effectively harnesses the strengths of both AdaBoost and Random Forest. By incorporating error correction from these models and enhancing robustness through a logistic regression meta-classifier, this approach optimizes the integration of diverse feature sets. The result is a significantly improved engagement prediction accuracy, reaching 63.09%, which outperforms all baselines. This superior performance underscores the efficacy of combining multiple learning strategies to refine predictive accuracy in complex datasets.
@@
Proposed V+P Late Fusion 63.09%
IV-D2 Choice of Machine Learning Models
Tables III and IV outline the machine learning models selected for visual and physiological modalities. The AdaBoost Classifier, achieving an accuracy of 43.03% for visual features, is chosen for its ability to enhance performance by focusing on difficult cases in complex visual data, such as variations in lighting and expression. In contrast, the Random Forest Classifier, which attains a higher accuracy of 55.25% for physiological features, is utilized due to its robustness in handling multivariate data and the feature selection capabilities essential for physiological signals such as heart rate and blood pressure peaks.
TABLE III: Choice of ML Models for Visual Modality
Model Accuracy
K-Nearest Neighbors Classifier 39.66%
Gradient Boost Classifier 42.77%
Support Vector Classifier 42.77%
Random Forest Classifier 42.86%
Decision Tree Classifier 42.95%
Ada Boost Classifier 43.03%
TABLE IV: Choice of ML Models for Physiological Modality.
Model Accuracy
Ada Boost Classifier 44.54%
Support Vector Classifier 46.76%
Decision Tree Classifier 51.67%
K-Nearest Neighbors Classifier 51.91%
Gradient Boost Classifier 54.48%
Random Forest 55.25%
IV-D3 Computational requirements analysis
Pre-extracting relevant features from raw video frames allows for parallel processing. Sequential task execution often results in idle processor time, reducing overall system efficiency. To address this, we implemented the ThreadPoolExecutor [45], which facilitates parallel processing by allowing multiple tasks to run concurrently across different threads. This approach not only optimizes CPU usage by reducing idle time but also enhances throughput. Table V details the performance improvements achieved with parallel processing techniques compared to traditional sequential methods, highlighting significant gains in processing speed and efficiency.
TABLE V: Comparison of Sequential (ST) and ThreadPoolExecutor (TPE) times in seconds.
@@ IV-D3 Computational requirements analysis
IV-E Discussion
The development of VisioPhysioENet introduces a novel approach in the area of multimodal engagement detection by leveraging both visual and physiological data streams. This innovative integration provides a more nuanced understanding of learner engagement, significantly improving both accuracy and robustness. With an accuracy of 63.09% on the DAiSEE dataset, VisioPhysioENet surpasses existing unimodal and multimodal methods. A distinct advantage of VisioPhysioENet lies in its advanced feature extraction techniques and efficient data integration strategy, which effectively handles real-world educational environments. The use of Dlib and OpenCV for visual feature extraction and the POS method for physiological signal analysis also ensures high performance with minimal computational overhead, making the system highly suitable for real-time applications that demand rapid and reliable engagement detection.
As shown in Tables III, II, and IV, the visual modality achieves 43.90% accuracy, while physiological signals alone yield 55.25%. By integrating these two streams, VisioPhysioENet attains 63.09%, illustrating how the multimodal approach captures complementary information often overlooked in unimodal systems. This combination leads to more robust performance and offers deeper insights into learners’ engagement states. Such synergy is particularly valuable in diverse educational settings where variations in gaze, facial expressions, and physiological responses contribute to the overall engagement profile.
Despite these strengths, the system faces challenges, especially regarding sensitivity to environmental variables such as lighting and motion artifacts. These factors can complicate data collection and introduce noise. Consequently, further refinement and additional testing are needed to ensure greater robustness across diverse populations and scenarios. Moreover, integrating physiological signals adds complexities in data capture and interpretation—small changes in emotional or physical state can alter physiological readings. Improved algorithms are crucial to distinguish between the variations that genuinely indicate engagement and those arising from unrelated factors.
VisioPhysioENet’s strong performance in utilizing multimodal (visual and physiological) data suggests its applicability in various applications. For example, in education, it can monitor student engagement levels, enhancing instructional strategies and ultimately improving learning outcomes [23]. In corporate training, this approach could strengthen employee learning retention [46]. Healthcare is another promising domain: by gauging patients’ understanding of treatment plans, clinicians can adjust patient education [2]. In advertisement, marketing, and consumer research, insights into consumer responses to media and advertising can be gleaned using this system [47]. Additionally, virtual events and conferences can benefit, as real-time feedback can inform dynamic content adjustments to heighten attendee engagement [42].
V Conclusion and Future Directions
VisioPhysioENet, a novel multimodal system that integrates visual cues and physiological signals, has been proposed to effectively detect learner engagement. Achieving a commendable accuracy of 63.09% on the DAiSEE dataset, it outperforms several state-of-the-art methods. This balance of accuracy and efficiency makes it a robust solution for real-world applications, especially suitable for dynamic environments.
Future enhancements will focus on improving scalability and robustness across diverse user groups and environments. This will include integrating additional modalities, such as audio cues and contextual data, to enrich the system’s understanding of learner engagement. Efforts will also aim to adapt the system for varied educational contexts, such as online learning platforms, virtual classrooms, and hybrid learning setups, ensuring its real-world adaptability and applicability in personalized learning experiences. Through these developments, VisioPhysioENet aims to become a versatile tool to advance intelligent educational technologies.